OpenStudy (anonymous):

multiply, simplify where possible. (sqrt8 - sqrt3)(sqrt8 +sqrt3)

jimthompson5910 (jim_thompson5910):

If you let x = sqrt(8) and y = sqrt(3), you get (x-y)(x+y) which becomes x^2-y^2

Parth (parthkohli):

Some trick here xD \(\Large \color{Black}{\Rightarrow (a - b)(a + b) = a^2 - b^2 }\)

OpenStudy (anonymous):

\[(\sqrt{8}-\sqrt{3})(\sqrt{8}+\sqrt{3})\]

jimthompson5910 (jim_thompson5910):

Then you can replace 'x' with sqrt(8) and y with sqrt(3) to get (sqrt(8))^2 - (sqrt(3))^2

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

(64-9)(64+9) ??

jimthompson5910 (jim_thompson5910):

not quite, I see what you're saying though

jimthompson5910 (jim_thompson5910):

(sqrt(8))^2 = 8 since the square and square root undo each other so (sqrt(3))^2 = 3 This means (sqrt(8))^2 - (sqrt(3))^2 becomes 8 - 3 = 5 So (sqrt(8) - sqrt(3))(sqrt(8) +sqrt(3)) = 5